What is 12X-10y=16. y = 0
What is 12X-10y=16. y = 0
Plug in y=0
12x-10(0)=16 any number multiply by 0 is zero
12x=16 divide both sides by 12
X=16/12 divisible by 4
X=4/3 ans

Answer by chiexpert(48)

What is 12X-10y=16. y = 0
Plug in y=0
12x-10(0)=16 any number multiply by 0 is zero
12x=16 divide both sides by 12
X=16/12 divisible by 4
X=4/3 ans